OpenStudy (anonymous):

Find an equation of the circle that satisfies the stated conditions. Center at the origin, passing through P(3, −6)

OpenStudy (anonymous):

Try x^2 + y^2= 45

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

The generic formula for a circle is: \[(X-Xo)^{2} + (Y-Yo)^{2} = r^{2}\] Where (Xo,Yo) is the center and r is the radius. In your case, you want it to be centered at the origin, so you make (Xo,Yo) = (0,0) That leaves you with: \[(X-0)^{2} + (Y-0)^{2} = r^{2}\] \[(X)^{2} + (Y)^{2} = r^{2}\] Now you have to pick a radius so that it crosses the point (3,-6).  You use Pythagoras and the h^2 is your radius. \[(X)^{2} + (Y)^{2} = (3)^{2}+(-6)^{2} = 9 + 36= 45\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

You're welcome. Hope that helped you understand. :)